OpenStudy (anonymous):

A cylinder has a radius of 12 m and a height of 9 m. What is the exact volume of the cylinder? 972pi m3 108pi m3 216pi m3 1296pi m3

Directrix (directrix):

Vcylinder = B*h where B is the area of the base which is a circle of radius 12 and h is the height. V = pi * 12^2 * 9 = ? @BrAnDoN_AAS

Directrix (directrix):

There's an online calculator here if you'd like to try it. https://www.google.com/?gws_rd=ssl#q=volume+of+cylinder

Directrix (directrix):

I approximated for pi and thought something was wrong with my work or the options. I do see the correct answer now. Do you?